40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Tue Apr 24 11:31:59 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling EX_fiber_MM.pdfEX_fiber_changes.png
    Reply  Tue Apr 24 18:54:15 2018, gautam, Configuration, ALS, Proposed changes to EX fiber coupling 
       Reply  Wed Apr 25 19:09:37 2018, gautam, Configuration, ALS, New look EX Fiber coupling EX_fiber_MM.pdfIMG_6977.JPGIMG_6972.JPGIMG_6974.JPGIMG_6976.JPG
          Reply  Thu Apr 26 11:24:50 2018, gautam, Configuration, ALS, New look EX Fiber coupling - pol stability polStab.pdf
             Reply  Thu Apr 26 18:58:21 2018, Bruce, Configuration, ALS, New look EX Fiber coupling - pol stability 
             Reply  Fri Apr 27 01:36:02 2018, gautam, Configuration, ALS, IR ALS noise performance 2018_04_BeatMouth_POX.pdf2018_04_BeatMouth.pdfALSSpecgram.pdfALS_ASD.pdfALSstab.pdf
                Reply  Wed May 2 21:39:33 2018, gautam, Configuration, ALS, IR ALS for EY 
                   Reply  Fri May 4 21:17:57 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table 
                      Reply  Tue May 8 00:40:51 2018, gautam, Configuration, ALS, BeathMouth pulled out of PSL table IMG_6986.JPGIMG_6987.JPG
                         Reply  Thu May 24 19:41:28 2018, gautam, Configuration, ALS, BeathMouth reinstalled on PSL table 
                      Reply  Thu May 24 13:06:17 2018, gautam, Configuration, ALS, DFD noises DFDcalib.pdfDFD_NB.pdf
                         Reply  Thu May 24 20:31:03 2018, gautam, Configuration, ALS, DFD noises TF_X_mag.pdfTF_X_phase.pdf
Message ID: 13786     Entry time: Tue Apr 24 18:54:15 2018     In reply to: 13784     Reply to this: 13789
Author: gautam 
Type: Configuration 
Category: ALS 
Subject: Proposed changes to EX fiber coupling 

I started working on the EX table. Work is ongoing so I will finish this up later in the evening, but in case anyone is wondering why there is no green light...

  1. EX laser shutter was closed.
  2. Disconnected EX input to the beat mouth at the PSL table in order to avoid accidentally frying the PDs.
  3. Prepared new optomechanics hardware
    • To my surprise, I found a bubble-wrapped K6XS mount (the one with locking screws for all DoFs) on the SP table. No idea where this came from or who brought it here, or how long it has been here, but I decided to use it nevertheless.
    • Prepared f = 200mm and f = -200mm lenses on traveling mounts (Thorlabs DT12, lenses are also Thorlabs, AR1064).
    • Made a slight translation of the beam path towards the north to facilitate going through the center of the mounted lenses.
    • Temporarily removed a beam dump from next to the final steering mirror before the Green REFL PD, and also removed one of the brackets between the enclosure and the table for ease of laying out components. These will be replaced later.
  4. Installed this hardware on the PSL table, roughly aligned beam path.
    • Beam now goes through the center of all lenses and is hitting the collimator roughly in the center.

To do in the eve:

  1. Clean fiber and connect it to the collimator.
  2. Optimize mode-matching as best as possible.
  3. Attenuate power using PBS and HWP so as to not damage the BeatMouth PD (Pthresh = 2mW). These are also required to make the polarizations of the EX coupled light (S-pol) and PSL (P-pol) go along the same axis of the PM fiber.
  4. Re-install temporarily removed beam dump and bracket on EX table.
  5. Re-install EX power monitoring PD.
  6. Measure beat frequency spectrum.
Quote:

Motivation: I want to make another measurement of the out-of-loop ALS beat noise, with improved MM into both the PSL and EX fibers and also better polarization control. For this, I want to make a few changes at the EX table. 

Barring objections, I will start working on these changes later today.


gautam 1245am: Fiber cleaning was done - I'll upload pics tomorrow, but it seems like the fiber was in need of a good cleaning. I did some initial mode-matching attempts, but peaked at 10% MM. Koji suggested not going for the final precisely tunable lens mounting solution while trying to perfect the MM. So I'll use easier to move mounts for the initial tuning and then swap out the DT12s once I have achieved good MM. Note that without any attenuation optics in place, 24.81mW of power is incident on the collimator. In order to facilitate easy debugging, I have connected the spare fiber from PSL to EX at the PSL table to the main EX fiber - this allows me to continuously monitor the power coupled into the fiber at the EX table while I tweak lens positions and alignment. After a bit of struggle, I noticed I had neglected a f=150mm lens in my earlier calculation - I've now included it again, and happily, there seems to be a solution which yields the theoretical 100% MM efficiency. I'll work on implementing this tomorrow..

ELOG V3.1.3-